# Balancing Intensity and Focality in Directional DBS Under Uncertainty: A Simulation Study of Electrode Optimization via a Metaheuristic L1L1 Approach

### v1

@@ Abstract
Background and Objective: As Deep Brain Stimulation (DBS) technology advances toward directional leads and optimization-based current steering, this study aims to improve the selection of electrode contact configurations using the recently developed L1-norm regularized L1-norm fitting (L1L1) method. The focus is in particular on L1L1’s capability to incorporate a priori lead field uncertainty, offering a potential advantage over conventional approaches that do not account for such variability.
@@
Deep Brain Stimulation (DBS) , Directional Lead , Anterior Nucleus of Thalamus , Convex Optimization , Metaheuristics
@@
[inst1]Computing Sciences Unit, Faculty of Information Technology and Communication Sciences, Tampere University, Tampere, Finland
@@ 1 Introduction
Deep Brain Stimulation (DBS) [1, 2] is a neuromodulatory therapy for managing movement disorders such as Parkinson’s disease [3, 4, 5, 6], essential tremor [7], refractory epilepsy [8, 9, 10], and dystonia [11], particularly in patients with insufficient response to pharmacological treatments [12]. The therapy delivers electrical stimulation to specific subcortical brain regions through implanted electrodes. Among the various parameters that influence clinical outcomes, the configuration of the stimulating electrodes plays a critical role, yet remains insufficiently standardized and not fully understood.
The central idea in determining an electrode configuration relies on predicting and controlling the regions of axonal activation induced by the lead [13, 14]. The extent of this activation depend on several factors, including tissue conductivity [15, 16], electrode geometry [17, 18], and the anatomical features of the surrounding brain structures [19]. Electrode configurations are typically classified by the number and polarity of active contacts within the stimulation target. In a monopolar (or unipolar) configuration, one or more cathodic contacts reside within the target structure, while the return path is an anode located at the implantable pulse generator, which is electrically distant and treated as a reference at infinity. This results in a broadly spherical electric field centered around the active contact(s). In contrast, bipolar configurations employ both cathodic and anodic contacts located within the brain, allowing for more localized current steering between adjacent contacts. Multipolar configurations generalize this further by utilizing multiple active contacts of both polarities within the lead, enabling finer control of the field shape and distribution.
Through deliberate selection of electrode configurations, clinicians can control which anatomical regions around the directional DBS (dDBS) leads are stimulated [20, 21, 18, 22, 23, 24, 25, 26]. This capability enables personalized shaping of the electric field in terms of spatial extent, intensity, and directionality, optimizing therapeutic outcomes while reducing off-target effects. However, the increasing complexity of modern lead designs has introduced a combinatorial explosion in possible configurations. This has driven significant research efforts into advanced electrode geometries and computational optimization strategies for electric field shaping [27, 28, 29, 30, 21, 31].
In this study, we focus on computational modeling of brain anatomy and optimization techniques to identify effective bipolar and multipolar electrode configurations for dDBS leads. We employ the L1-norm regularized L1-norm fitting (L1L1) method [32], a sparse optimization framework designed to yield focal stimulation profiles by simultaneously penalizing the control vector and fitting residuals using L1-norms. This formulation promotes sparsity in the solution space, enabling interpretable and spatially selective activation patterns. Moreover, the L1L1 method naturally incorporates uncertainty in the lead field model, increasing robustness to measurement noise and reducing the risk of overfitting. These properties make it well-suited for individualized DBS modeling, where anatomical variability and signal noise present significant computational and neurophysiological challenges.
To constrain the optimization problem, we define the fitting region based on the spatial attenuation of the lead field, which approximates the clinically relevant volume of tissue activated (VTA). This region corresponds to the subset of neural tissue where the electric field amplitude exceeds a threshold sufficient for neuronal activation. Regions outside this boundary are excluded from the fitting procedure, as stimulation effects there are negligible or uncertain. Employing the VTA as a constraint provides a computationally tractable surrogate for detailed multicompartment neuron models that simulate extracellular potentials and neuronal activation dynamics, which are often prohibitively expensive in high-dimensional optimization tasks [33, 34, 35].
To quantitatively assess the robustness of the L1L1 optimization framework under realistic conditions, we expand the study to introduce additive Gaussian noise to the lead field matrix, simulating possible modeling errors and measurement variability. The noise magnitude is calibrated to reflect plausible physiological and instrumentation uncertainties such as strong deformation/distortions. We then evaluate the L1L1 method against two established baseline methods: the Reciprocity Principle (RP) [36], which leverages reciprocity theory to infer optimal stimulation patterns, and Tikhonov Regularized Least Squares (TLS) [37, 38], which balances data fidelity and solution smoothness through L2-norm regularization. This comparative analysis elucidates the trade-offs in spatial precision, sparsity, and noise resilience among the methods, highlighting the advantages of the L1L1 approach for individualized electrode configuration optimization.
@@ 2 Materials and Methods
2.1 Head Model Discretization
This study employs finite elements (FE) to solve the linear forward problem derived from the complete electrode model [39] and to illustrate the interpolated distributions on the FE mesh [40, 41]. The FE mesh is constructed using an openly accessible T1-weighted MRI-based dataset [42], which includes MRI data from a healthy 48-year-old right-handed individual who serves as our representative subject. The brain volume is segmented using the FreeSurfer Software Suite [43], and subcortical structures are extracted from FreeSurfer’s Aseg atlas (Fig. 1(a)). The head model’s parcellation is based on a multicompartment conductivity distribution, comprising 18 cortical and subcortical segments. The element-wise constant electrical conductivity values for these regions were set according to [44], and the subcortical nuclei values were taken from [45, 46]. Using this curated data, an unstructured tetrahedral boundary-fitted FE mesh is generated using the Zeffiro Interface (ZI) [47], with a resolution of 3.0 mm (millimeters). The volume of the thalamus compartment was further refined through three nested refinements [48] with each tetrahedron subdivided into four congruent parts, resulting in tetrahedra with diameters ranging from 0.1725 mm to 0.375 mm. This process led to a mesh consisting of 4.4 million nodes and 25.4 million elements (Fig. 1(b)). We define the anterior nucleus of the thalamus (ANT) as the stimulation target, similar to ANT-DBS, a clinically established approach for treating epilepsy [8, 9].
@@ 2.1 Head Model Discretization
(b) Refined and boundary-fitted thalamus sub-compartment.
@@
Figure 1: (1(a)): A head model including skull, forebrain, midbrain, hindbrain structures as well as directional deep brain stimulation (dDBS) lead positioned at the right-hemispheric anterior nucleus of thalamus (ANT). (1(b)): Refined and boundary-fitted thalamus sub-compartment with the distribution of tetrahedral mesh edges and an interpolated current intensity distribution shown on the surface. (1(c)): Placement of a 40-contact dDBS lead with a bipolar electrode configuration (anodal as red, and cathodal as blue). The dipolar target current (red) is depicted by the red pin, indicating the target position and orientation; a 6.0 millimeter (mm) region of interest (ROI) is displayed as an arrangement of source points (yellow squares) for High-Resolution Lead Field (HR-LF) matrix. (1(d)): Probing strategy of the 8- and 40-electrode contacts inspired by Abbott’s St. Jude Medical Infinity™ Directional Lead 6180 model and Medtronic-Sapiens, respectively.
2.2 Lead Field Matrix
The linear forward mapping can be formulated as a real $N\times K$ lead field matrix $\hat{\bf L}$, which relates current injections at the electrodes to the resulting volumetric current density distribution across different electrode configurations. This matrix characterizes the numerical relationship between the real $K\times 1$ current pattern ${\bf y=(y_{1},\ldots,y_{K})}$ and the real $N\times 1$ discretized volume current density $\hat{\bf x}=(\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{N})$. The latter variable represents an arrangement of uniformly distributed source points within the target domain, each with three degrees of freedom (DOFs) corresponding to three dipolar unit currents with Cartesian orientations.
We generate a lead field matrix for two different resolution levels: (i) Low-Resolution Lead Field (LR-LF), which corresponds to a target set of 244 $\times$ 3 spatial DOFs that cover the entire thalamus compartment. This robust interpretation offers a functional assessment of the model, serving as a benchmark for our numerical experiments. To validate whether our minimal interpretation is adequate, we compare the volumetric current density distribution with a locally detailed lead field matrix, (ii) High-Resolution Lead Field (HR-LF). This matrix consists of 3620 $\times$ 3 DOFs distributed across the thalamus, meaning that the mutual distance between nearest neighboring DOFs is reduced by 40% compared to LR-LF. Due to the increased computational cost of using a larger lead field size, our examination of HR-LF is confined to a spherical region of interest with a 6.0 mm radius, co-centered with the lead model (Fig. 1(c)).
2.3 Probing Strategy
@@
The maximum total current injection supplied by all the active electrodes is limited to 4.0 mA (milliamperes) (5), and the maximum current per electrode contact is set to $\pm$ 2.0 mA (6). This can be considered a moderately typical setting, for example, 3.0 mA as a reference for monopolar patterns in [50]. The total absolute sum of electric potential corresponding to the contacts in the lead is normalized to meet the intended maximum current injection value (7). As a reference threshold for neural activation, we consider the current amplitude of 3.85 A/m2 (amperes per square meter) [51], which falls within the excitation current thresholds of 2.5 to 6.0 A/m2 estimated for nerve fibers in the upper limb area of the motor cortex at frequencies of 50.0 Hz and 2.44 kHz [52]. Our optimization framework maintains this reference threshold constant across all tested configurations. By holding temporal characteristics constant, we ensure that the observed variations in activation patterns arise from differences in spatial registrations and current steering, rather than from stimulation timing. This choice follows the framework established in previous studies by Deli et al. [53] and Schlaepfer et al. [54], where monopolar and bipolar stimulation configurations were evaluated using a consistent stimulation protocol, specifically maintaining a pulse width of 60 µs and a frequency of 130 Hz across all conditions. This design allowed the authors to isolate the effects attributable solely to electrode configuration by holding other key stimulation parameters constant, with amplitude as the only adjusted variable.
2.4 Optimization Algorithms
The aim of this study is the application of mathematical optimization algorithms to retrieve the current pattern ${\bf y}$ (4) within the local target region, aiming to optimally match the volumetric current distribution generated by the target current dipole modeling [55, 56, 51]. To quantify the optimal electrode configurations derived from the algorithm, we compare them based on the decision variables $\Gamma$, $\Xi$, and $\Theta$, as delineated by the optimization paradigm. The focused current density $\Gamma$ (A/m2) is defined as
${\Gamma}=\frac{{\bf x}_{1}^{T}{\bf L}_{1}{\bf y}}{\|{\bf x}_{1}\|_{2}}.$
The focality of the stimulation, or field ratio, $\Theta$, is defined as the ratio of the focused current density divided by the undesired nuisance current density ($\Xi$) and is defined as
${\Theta}=\frac{{\Gamma}}{\Xi}\quad\hbox{with}\quad\Xi=\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}.$
@@ 2.4.1 Bipolar Configurations
Bipolar configurations are derived through the reciprocity principle (RP) [36]. The principle states that the maximum focused current density, i.e., $\Gamma_{\hbox{\scriptsize{max}}}$, of an applied linear system can be achieved by limiting the electrode configuration to a pair of anodal and cathodal electrode contacts (Fig. 3). This configuration is obtained by multiplying the dipolar target current distribution by the transpose of the lead field matrix, selecting both the greatest positive and negative poles, and setting the remaining electrodes in the lead to zero. The reciprocity principle aims at maximizing current density at the expense of field ratio fidelity.
@@ 2.4.2 Multipolar Configurations
Multipolar configurations, i.e., leads with multiple active electrode contacts, are obtained via convex optimization techniques such as L1-norm regularized L1-norm fitting (L1L1) [32] and through Tikhonov-Regularized Least Squares (TLS) [37, 38] methods. Both methods employ a metaheuristic search strategy to explore only a carefully chosen fraction of an otherwise huge search space—trading guaranteed optimality for scalability and adaptability—to handle large, ill-behaved, or poorly understood problems within practical time and memory constraints [57]. The candidate solution is the one that can achieve the highest ratio between the (target) current density at the stimulation site and that in the surrounding adjacent areas (referred here to as nuisance current). The determination of this solution is defined by tuning a set of predefined hyperparameters along with the current pattern ${\bf y}$, subject to specific metacriteria, namely, maximizing the field ratio while ensuring $\Gamma\geq\Gamma_{0}>0$, thereby maintaining sufficient intensity at the target location (with $\Gamma_{0}=0.80$ mA in this case). Without a lower bound for intensity $\Gamma$, the intensity of the $\Theta$-maximizer is likely to diminish. A rough upper bound is provided as a L1-norm fitting threshold to suppress nuisance current density $\Xi$. In contrast, traditional least-squares formulations tend to produce smoothly distributed currents, which may be clinically suboptimal, as focal and sparse activation is often preferred to minimize side effects and enhance therapeutic precision. Moreover, least-squares approaches are sensitive to noise and outliers, often yielding non-robust solutions [58]. Detailed formulation of both the L1L1 and TLS methods is provided in A.

### v2

# Balancing Intensity and Focality in Directional DBS Under Uncertainty: A Simulation Study of Electrode Optimization via Metaheuristic L1L1 Method

@@ Abstract
Background and Objective: As Deep Brain Stimulation (DBS) technology advances toward directional leads and optimization-based current steering, this study aims to improve the selection of electrode contact configurations using the recently developed metaheuristic L1-norm regularized L1-norm fitting (L1L1) method. The focus is in particular on L1L1’s capability to incorporate a priori lead field uncertainty, offering a potential advantage over conventional approaches that do not account for such variability.
@@
Deep Brain Stimulation (DBS) , Anterior Nucleus of Thalamus , Convex Optimization , Metaheuristics
@@
[…]
@@ 1 Introduction
Deep Brain Stimulation (DBS) [volkmann2002introduction, Perlmutter_2006_DBS, Gardner_2013_history_DBS] is a neuromodulatory therapy used to alleviate symptoms in patients with movement and neurological disorders, including, e.g., Parkinson’s disease [rodriguez2005bilateral, picillo2016programming, krack2019deep, steffen2020bipolar], essential tremor [benabid1996chronic], refractory epilepsy [LEHTIMAKI_2016_DBS_targetting, Jarvenpaa_2020_Improving_ANTDBS, fasano2021experience], and dystonia [johnson2013neuromodulation], by implanting an electrode lead into a targeted subcortical region and delivering controlled electrical stimulation through its contacts to modulate pathological neural activity. The central objective when determining this region of axonal activation is to shape the resulting volume of tissue activated (VTA), elicited by activating specific electrode contacts along the lead, so that it aligns with the intended therapeutic target [mcintyre_2004_cellular, butson_2006_electrodesign, butson_2007_patient, miocinovic_2006_subthalamic].
Computational representations of DBS activation patterns are often categorized by the number and polarity of active contacts. In monopolar configurations, one or more cathodic contacts are driven against the distant implantable pulse generator (IPG), which serves as a reference at infinity [volkmann_basic_algorithm_DBS_2006]. Bipolar configurations confine both anode and cathode to contacts on the lead, creating a locally circulating current and producing a more spatially restricted field [steffen2020bipolar, DELI_2011_UnipolarBipolarComparison]. Multipolar configurations further extend the programmable space by allowing multiple contacts with mixed polarities to shape the field in a more flexible manner. As contemporary leads incorporate segmented or high-density arrays, the resulting combinatorial growth in possible patterns renders exhaustive manual programming infeasible [Anderson_2018].
Effective steering of the VTA depends on a combination of tissue properties [butson_2005_tissue, McIntyre_2006_OptimizingDBS], electrode geometry and orientation [willsie2015computational], and the anatomical variability of the target region [wei_2005_currentdensityDBS, van2015novel]. Clinical investigations such as [contarino2014directional, pollo2014directional, dembek2017directional] and the synthesis in [Steigerwald_dDBS_2019] demonstrated directional steering empirically through stimulation mapping. Complementary computational studies, including [alonso2016investigation, zhang2020steering], examined how geometry, contact orientation, and heterogeneous tissue conductivities influence field distribution and VTA through finite element modeling. Although [Anderson_2018] illustrates how convex formulations can be incorporated into directional lead control, it represents one contribution within a broader set of emerging model-based strategies aimed at systematically optimizing stimulation patterns for directional DBS.
We employ the metaheuristic L1-norm regularized L1-norm fitting (L1L1) method [galazprieto_2022_L1vsL2] to retrieve electrode configurations over standard 8-contact and a complex 40-contacts directional DBS leads. The L1L1 method treats both the control vector and the data misfit with L1 penalties, which makes the optimization resistant to noise and model mismatch while discouraging diffuse or overly complex solutions, unlike least-squares formulations. We include configurations following Reciprocity Principle (RP2) [FERNANDEZCORAZZA2020116403], which leverages reciprocity theory to infer optimal stimulation patterns, and Tikhonov Regularized Least Squares (TLS) [dmochowski2011optimized], which balances data fidelity and solution smoothness through L2-norm regularization. For clarification: the terms “L1” and “L2” refer to the discrete 1- and 2-norms, respectively, of finite-dimensional vectors. The capitalized form L1L1 is used solely as the algorithm name and does not refer to the Lebesgue functional spaces, i.e., $L^{1}$ or $L^{2}$.
@@ 2 Materials and Methods
The quasi-static forward problem was modeled using the Complete Electrode Model (CEM) formulation [pursiainen2017forward] and solved with the Finite Element Method (FEM). Electrical field distributions were interpolated over the volumetric mesh following [pursiainen2016, miinalainen2019]. All modeling, meshing, and optimization procedures were performed using the Zeffiro Interface (ZI) version 6.1.0 [pursiainen2025zeffiro], an open-source MATLAB-based pipeline that provides unstructured boundary-fitted, multi-compartment tetrahedral FE mesh generation from MRI-derived segmentations.
2.1 Realistic Head Model
We generated a tetrahedral-based mesh with a nominal element size of 3.0 mm from an open access T1 weighted MRI data set from a neurologically healthy adult [piastra_maria_carla_2020_3888381]. We performed cortical segmentations using FreeSurfer [fischl2012freesurfer], supported by functions from SPM12 package [ASHBURNER20091163]. Deep nuclei labels were extracted according to the Aseg atlas (Fig. 1(a)) resulting in a realistic head model with 18 anatomically distinct tissue compartments. A double-layered cylindrical compartment with a 1.27 mm outer diameter, and a 0.5 mm thick encapsulation layer [anderson_2019_anodic] was further added and interpolated using in-house scripts. The cylinder functions as a proxy for setting the electrodes with their respective geometric properties.
As our target sub-compartment would be then the anterior nucleus of the thalamus (ANT), consistent with ANT-DBS applications in epilepsy [LEHTIMAKI_2016_DBS_targetting, Jarvenpaa_2020_Improving_ANTDBS], the thalamus compartment underwent three levels of nested refinements subdividing every tetrahedron on each step into four congruent sub-elements [galazprieto_2023_mesh]. This procedure yielded element diameters ranging from 0.1725 to 0.375 mm in the refined region and produced a final mesh comprising approximately 4.4 million nodes and 25.4 million tetrahedral elements (Fig. 1(b)).
We assigned electrical conductivities as follows: 0.14 S/m for white matter, 0.33 S/m for gray matter, 0.0064 S/m for skull, 1.79 S/m for cerebrospinal fluid (CSF), 0.33 S/m for subcortical structures, and 0.33 S/m for skin, as reported in [dannhauer2011modeling]. Subcortical nuclei were assumed to share electrical conductivity values of gray matter, while ventricular regions were assigned the conductivity of CSF, in accordance with [rezaei2021reconstructing] and its referenced sources.
2.2 Lead Field Matrix
To establish a baseline representation for spatial attenuation of the lead field and its correspondence to feasible VTA coverage, we derived an initial low-resolution lead field (LR-LF) comprising $244\times 3$ dipolar DOFs distributed throughout the thalamus. The spatial extent for detailed current steering was then guided by reported activation radii around the electrodes, where typical stimulation produces tissue activation on the order of approximately 2.0–4.0 mm [butson_2007_patient, miocinovic_2006_subthalamic], and may extend to roughly 3.8–4.0 mm under clinically relevant parameters [madler2012activated]. We then establish a spherical region of 6.0 mm radius centered on the lead to represent the domain in which fine-grained modulation is physiologically meaningful. A high-resolution lead field (HR-LF) was subsequently generated by increasing the dipole density to $3620\times 3$ DOFs across the same anatomical region and then retaining only those sources falling within the 6.0 mm domain (Fig. 1(c)). A full derivation can be found in A.1.
@@ 2.2 Lead Field Matrix
(b) Refined unstructured boundary-fitted thalamus sub-compartment with activity (smoothed).
@@
Figure 1: (1(a)): A head model including skull, forebrain, midbrain, hindbrain structures as well as a directional deep brain stimulation (DBS) lead positioned at the right-hemispheric anterior nucleus of thalamus (ANT). (1(b)): Refined unstructured boundary-fitted thalamus sub-compartment with the distribution of tetrahedral mesh edges and an interpolated current intensity distribution shown on the surface. (1(c)): Placement of a 40-contact dDBS lead with a bipolar electrode configuration (anodal as red, and cathodal as blue). The dipolar target current (red) is depicted by the red pin, indicating the target position and orientation; a 6.0 millimeter ($m$) fitting region is displayed as an arrangement of source points (yellow squares) for High-Resolution Lead Field (HR-LF) matrix. (1(d)): Probing strategy of the 8- and 40-electrode contacts inspired by Abbott Infinity™ and Medtronic-Sapiens, respectively.
2.3 Optimization Framework
The applied optimization framework aims to return a focal anodal–cathodal electrode configuration that best reproduces the volumetric current distribution associated with a prescribed target current dipole model [demunck1988, mosher2002multiple, medani_2012_ModellingDipoles]. The framework enhances field components aligned with the target while attenuating directions dominated by nuisance activation. Active electrodes are selected iteratively by adjusting the current pattern $\mathbf{y}$ under a set of predefined hyperparameters and an expected number of active contacts $K$, depending on the chosen algorithm (e.g., Reciprocity Principle, L1L1 Method, or Tikhonov-regularized least squares). Metaheuristics are employed to explore the feasible design space and identify current patterns that best satisfy the required performance characteristics [galazprieto_2024_lattice].
To restrict the number of admissible solutions, we impose a lower bound on the focused current density $\Gamma$ ($\Gamma\geq\Gamma_{0}=0.80~\mathrm{mA}$) to ensure sufficient activation at the target, as well as an upper bound on the nuisance current density $\Xi$ to limit excessive off-target stimulation. The candidate solution is thus defined as the current pattern that maximizes the field ratio between target- and nuisance-region current densities while satisfying these constraints.
As safety constraints, the total (absolute) injected current is constrained to $\mu=4.0~\mathrm{mA}$ (6), with per-contact limits $\gamma=\pm 2.0~\mathrm{mA}$ (7). Electrode potentials are normalized to satisfy the net-current constraint (8). Lastly, the stimulation is, however, treated as time-invariant, as the objective of this work is to characterize the spatial organization of current flow and the resulting current density distributions, rather than to model pulse dynamics or frequency-dependent neuronal responses.
For reference magnitude $\|{\bf x}_{1}\|_{2}=3.85$ A/m2 corresponds to the activation threshold reference. This value is obtained by dividing the average dipole moment density of pyramidal neurons (0.77 nAm/mm2) by an approximate dendritic axis length of 0.2 mm [MURAKAMI201549].
2.3.1 Electrode Configurations
@@ 2.3.2 Bipolar Configurations
Bipolar patterns are obtained using the reciprocity principle (RP) [FERNANDEZCORAZZA2020116403]. The principle states that the maximal focused current density $\Gamma_{\mathrm{max}}$ for a linear quasi-static system is achieved by selecting a single anode–cathode pair. In practice, the target dipolar field is multiplied by the transpose of the lead field matrix, and the contacts corresponding to the strongest positive and negative entries are chosen as anode and cathode, respectively, with all remaining contacts set to zero (Fig. 2). RP maximizes the delivered current density at the target but does not optimize the ratio between target activation and nuisance activation, and thus may yield suboptimal field selectivity.
@@ 2.3.3 Multipolar Configurations
Multipolar configurations are derived using the L1-norm regularized L1-norm fitting (L1L1) method [galazprieto_2022_L1vsL2] and Tikhonov-regularized least squares (TLS) [dmochowski2011optimized]. Both methods assume a predetermined set of candidate active contacts and estimate the corresponding stimulation amplitudes under polarity, charge, and safety constraints (detailed in A.2).
The L1L1 formulation imposes an L1 penalty on both the stimulation vector and the residual, which enhances robustness to lead-field uncertainty and promotes sparse, interpretable support. TLS, in contrast, employs an L2 penalty that stabilizes the solution in the presence of correlated columns and numerical ill-conditioning. Because the full combinatorial search space of possible contact subsets grows exponentially with $K$, both formulations employ a metaheuristic support-selection strategy to efficiently explore only a tractable subset of candidate configurations while maintaining practical computational cost [galazprieto_2024_lattice].
8-contact Electrode Configuration
40-contact Electrode Configuration
Figure 2: Examples of current injection patterns obtained using the Reciprocity Principle (RP), metaheuristic L1-norm regularized L1-norm fitting (L1L1) method under two different nuisance current thresholds (L1L1(A) with $\varepsilon=\{0:-160\}$ and L1L1(B) with $\varepsilon=\{0:-10\}$), and the Tikhonov-Regularized Least Square (TLS) method. The target position and orientation (parallel to the lead) is indicated by the red upward pin. The bar plots display the resulting electrode current amplitudes, while the contact maps visualize their spatial distribution over the electrodes; cathodal (blue), anodal (red), and inactive (gray/green). The RP method yields a bipolar configuration maximizing focused current density $\Gamma$, while the remaining methods searches for a multipolar configuration maximizing focality $\Theta=\Gamma/\Xi$ under an a priori constraint $\varepsilon$ on the nuisance current density $\Xi$, determined by the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ and the lead field attenuation. The L1L1(B) result reflects more strongly constrained $\Xi$ (smaller dynamic range), activating fewer contacts whereas L1L1(A) and TLS allow higher $\Theta$ values and broader excitation patterns. The voltage/current on the electrode contacts are $\pm 2.0$ mA (milliamperes).